OpenStudy (anonymous):

Solve the system of equations below by graphing them with a pencil and paper. Enter your answer as an ordered pair y = x + 5 y = -2x - 1

OpenStudy (anonymous):

x+5=-2x-1 3x=-6 x=-2 sub to y=x+5 y=3 (-2,3)

OpenStudy (anonymous):

Just out of boredness, I'll solve this by substitution as well. y = x + 5 MULTIPLY by 2 y = -2x - 1 2y = 2x + 10 y = -2x - 1 ____________ 3y = 9 3y/3 = 9/3 y = 3 Sub Y into the either equation. (Since it was solved via first equation, I'll do the second one) (3) = -2x - 1 3 + 1 = -2x 4/-2 = -2x/-2 -2 = x (x,y) (-2,3)